A bullet of mass 0.02kg travelling with speed 100m/s comes to rest when it has gone 0.4m into sand. Find the resisting force exerted by the sand
1
Expert's answer
2020-03-12T02:50:09-0400

"A=Wk1-Wk2=Wk1; 
A=F \times S=Wk1=\frac{m\times v^{2}}{2};
F=\frac{m\times v^{2}}{2\times S}=250 N"